OpenStudy (anonymous):

Find the value of the 150th derivative of e^(x^50) at x = 0. I think I should use a Maclaurin series but I don't know how. Thanks

OpenStudy (anonymous):

\[e^u=\sum_0\frac 1{n!}u^n\] \[e^{x^a}=\sum_0\frac 1{n!}x^{an}\]

OpenStudy (anonymous):

\[\sum_0\frac 1{n!}x^{an}\] \[\sum_1\frac 1{n!}~an~x^{an-1}\] \[\sum_2\frac 1{n!}~an(an-1)~x^{an-2}\] ... \[\sum_{150}\frac 1{n!}~an(an-1)...(an-149)~x^{an-150}\]

OpenStudy (anonymous):

we can adjust the index now to 0 by -150 +150 \[\sum_{150-150}\frac 1{(n+150)!}~a(n+150)(a(n+150)-1)...(a(n+150)-149)~x^{a(n+150)-150}\]

OpenStudy (anonymous):

almost forgot that a = 150 too :)

OpenStudy (anonymous):

er ..a = 50 that is

OpenStudy (anonymous):

It should somehow be 150! / 6, I can't see how.

OpenStudy (anonymous):

well, at x=0, everything but the first term goes zero, so lets test out my idea

OpenStudy (anonymous):

\[\frac 1{(150)!}~50(150)(50(150)-1)...(50(150)-149)\] the wolf might work for a quick determination :)

OpenStudy (anonymous):

Well wolfram agrees but this was an old exam question so there must be a simpler way.

OpenStudy (anonymous):

\[\frac 1{(150)!}~k(k-1)(k-2)...(k-149)\] k=50(150) = 7500

OpenStudy (anonymous):

yeah, i think im off in my coding .... but that is the process i would attempt

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=150!%2F6%2C+1%2F150!+*+Product [7500-x%2C+{x%2C0%2C149}] Seems to be a bit off, but I like the approach style!

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

0: e^u 1: u' e^(u) 2: u'' e^(u) + (u')^2 e^(u) 3: u''' e^(u) + 2(u')u'' e^(u) + u'' u' e^(u) + (u')^3 e^(u) hmmm this idea doesnt seem to be evening out too well ... mac series is the way to go fer sure

OpenStudy (anonymous):

Seems to be that way! I'll keep the question open to see if someone else has an idea! Thank you for all your help!

OpenStudy (anonymous):

youre welcome, and good luck :)